OpenStudy (anonymous):

Two angles that form a linear pair have a ratio of 2 : 3. What is the measure of both angles?

OpenStudy (anonymous):

Linear pair? Does that mean that both angles add up to 180?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

2x+3x=180 Solve for x

OpenStudy (anonymous):

x = 36

OpenStudy (anonymous):

2x+3x=180 x(2+3)=180 x(5) = 180 x =180/5= 36 First angle is 36*2 Second angle is 36*3

OpenStudy (anonymous):

72 & 108 yes?

OpenStudy (anonymous):

Yea that's what I got. If your definition of linear pair means that both angles add up to 180.

OpenStudy (anonymous):

Thank you!